Differential Analysis for a Discontinued Product

A condensed income statement by product line for Healthy Beverage Inc. indicated the following for Fruit Cola for the past year:

Sales $234,900
Cost of goods sold 109,000
Gross profit $125,900
Operating expenses 143,000
Loss from operations $(17,100)

It is estimated that 15% of the cost of goods sold represents fixed factory overhead costs and that 21% of the operating expenses are fixed. Because Fruit Cola is only one of many products, the fixed costs will not be materially affected if the product is discontinued.

a. Prepare a differential analysis dated January 5 to determine whether Fruit Cola should be continued (Alternative 1) or discontinued (Alternative 2). If an amount is zero, enter "0". Use a minus sign to indicate a loss.

Differential Analysis
Continue Fruit Cola (Alt. 1) or Discontinue Fruit Cola (Alt. 2)
January 5
Continue Fruit Cola (Alternative 1) Discontinue Fruit Cola (Alternative 2) Differential Effect on Income (Alternative 2)
Revenues $ $ $
Costs:
Variable cost of goods sold
Variable operating expenses
Fixed costs
Income (Loss) $ $ $

b. Should Fruit Cola be retained?
Answer #1
a
Continue Cups (Alternative 1) Discontinue Cups (Alternative 2) Differential Effect on Income (Alternative 2)
Revenues 234900 0 -234900
Costs:
Variable cost of goods sold -92650 0 92650
Variable selling and admin. expenses -112970 0 112970
Fixed costs -46380 -46380 0
Income (Loss) -17100 -46380 -29280
b
Yes, Fruit Cola should be retained.
If Fruit Cola is Discontinued, net income will decrease by $29280
Workings:
Variable cost of goods sold =109000*(1-15%)
Variable selling and admin. expenses =143000*(1-21%)
Fixed costs =(109000*15%)+(143000*21%)